. If 10.0g of an organic compound is combusted in excess O2(g) to give 20.0g of Carbon (IV) oxide and 5.0g of water vapour, what is the molecular formula of the compound? Take molar mass of the compound to be 83.947g/mol.


1
Expert's answer
2021-08-28T06:27:10-0400

Let the organic compound be X

Moles of X"=\frac{10}{83.947}=0.12"



Moles of "Co_2 =\frac{20.0}{44.01}=0.45"



Moles of "H_2O=\frac{5}{18.01}=0.28"


Ratios "\frac{0.12}{0.12}=1,
\frac{0.45}{0.12}=3.75=4, 
\frac{0.28}{0.12}=2.33=2"



Formula "X(CO_2)_4(H_2O)_2"